If the hypotenuse of a right triangle is 15.3m and the angle are 41 degrees and 49 degrees, what is the length of the leg opposite the 49 degree angle?


1
Expert's answer
2021-02-03T16:12:44-0500

By definition, the sine of an angle in a right triangle is:


"\sin\alpha = \dfrac{\text{opposite leg}}{\text{hypotenuse}}"

Taking "\alpha = 49\degree", and "\text{hypotenuse} = 15.3m" obtain the opposite leg:


"\text{opposite leg} = 15.3m\cdot \sin{49\degree} \approx 11.55m"

Answer. 11.55 m.